Circular No. 8837

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2007 JA_{21} (LINEAR) An apparently asteroidal object reported by the LINEAR survey (dis-

covery observation posted below; prematurely announced as 2007 JA_{21} on MPEC 2007-J52) has been found by several observers to show cometary appearance following posting on the Minor Planet Center’s ‘NEOCP’ webpage. Q.-z. Ye reports that three stacked 30-s CCD exposures obtained with the Lulin Observatory 0.41-m Ritchey-Chrétien reflector on May 12.82 UT show a coma of diameter ∼ 3′′ with an apparent but uncertain very faint tail around 7′′-10′′ long in p.a. ∼ 45^{o}. J. G. Ries writes that her Rband images obtained with the McDonald Observatory 0.76-m reflector on May 13.4 show the comet to be diffuse with a tail ≈ 10′′–13′′ long toward the north-northeast. Also, G. Sostero reports that forty co-added 120-s unfiltered CCD frames taken by L. Donato, M. Gonano, V. Gonano, E. Guido, and himself with a 0.45-m reflector at Remanzacco, Italy, during May 13.91–13.98 show a compact coma nearly 10′′ in diameter (giving the object’s magnitude as 18.7–18.8).

2007 UT  α_{2000}  δ_{2000}  Mag.
May 11.30048  17^{h}40^{m}15.^{s}69  +53^{o}06′22′′.3  19.9

Additional astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2007-J55:

T = 2007 May 11.272 TT  ω = 113.^{o}384
Ω = 66.274 }2000.0
q = 5.45549 AU  i = 89.472

COMET P/2007 H1 (McNAUGHT) Additional observations of this comet (cf. IAUC 8830), published on

MPEC 2007-J35, yield the following elliptical orbital elements:

T = 2007 Aug. 8.740 TT  ω = 199.^{o}594
e = 0.42222  Ω = 144.670 }2000.0
q = 2.31257 AU  i = 11.802
a = 4.00251 AU n^{o} = 0.123085 P = 8.0 years
2007 May 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green